OpenStudy (anonymous):

Please explain your answer and show work! -1+6+(-7)=?

jimthompson5910 (jim_thompson5910):

what is -1 + 6

jimthompson5910 (jim_thompson5910):

any ideas?

OpenStudy (anonymous):

@jim_thompson5910 - -5?

OpenStudy (anonymous):

-5

jimthompson5910 (jim_thompson5910):

think of it like this

jimthompson5910 (jim_thompson5910):

you are in debt 1 dollar (represented by -1) you then find 6 dollars somewhere how much money do you have overall now?

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

oh and once you find the 6 dollars you have to pay off that debt immediately

jimthompson5910 (jim_thompson5910):

yep so that means -1 + 6 = 5

OpenStudy (anonymous):

nice explaining jim :)

jimthompson5910 (jim_thompson5910):

so we go from -1+6+(-7) to 5 + (-7)

jimthompson5910 (jim_thompson5910):

thanks kelliegirl33, I really appreciate it

OpenStudy (anonymous):

-2

jimthompson5910 (jim_thompson5910):

to evaluate 5 + (-7), use the idea of money again you have 5 dollars, then you spend 7 dollars and you'll end up 2 dollars in debt so good, 5 + (-7) = -2

jimthompson5910 (jim_thompson5910):

overall, -1+6+(-7) = -2

OpenStudy (anonymous):

so if I had to show work on my paper how would it look?

jimthompson5910 (jim_thompson5910):

basically you could add the numbers like we did where we added them in pairs and we broke up the problem if you had more numbers, you just keep the same basic pattern going ex: 1+2+3+4 3+3+4 ... add the first pair 6+4 ... add the next pair 10 ... you keep adding pairs until you're done

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you so much!!! This was a HUGE help to me! :D Medals for you. :)

jimthompson5910 (jim_thompson5910):

you're welcome, glad to be of help